OpenStudy (anonymous):

write an equation in slope intercept form of the line through point P(6,7) with slope -2

OpenStudy (helder_edwin):

the formula of the slope is \[\large m=\frac{y_1-y_2}{x_1-x_2} \] u have m=-2 and (x1,y1)=(6,7). replace and do the algebra

OpenStudy (anonymous):

i just need to know the formula, i am still a little confused, can you explain it a little more? :)

OpenStudy (helder_edwin):

we have \[\large -2=\frac{7-y}{6-x} \] now do the algebra

OpenStudy (anonymous):

so is it y-7=-2(x-6) ?

OpenStudy (helder_edwin):

yes but u have to do some algebra \[\large y-7=-2(x-6)=-2x+12 \] \[\large y=-2x+19 \] which is in the slope-intercept form

OpenStudy (anonymous):

oh thank you!